# Understanding Practical Membership Privacy of Deep Learning

### v1

@@ Abstract
We apply a state-of-the-art membership inference attack (MIA) to systematically test the practical privacy vulnerability of fine-tuning large image classification models. We focus on understanding the properties of data sets and samples that make them vulnerable to membership inference. In terms of data set properties, we find a strong power law dependence between the number of examples per class in the data and the MIA vulnerability, as measured by true positive rate of the attack at a low false positive rate. For an individual sample, large gradients at the end of training are strongly correlated with MIA vulnerability.
@@ 1 Introduction
Machine learning models are prone to memorising their training data, which makes them vulnerable to privacy attacks such as membership inference attacks (MIAs; Shokri et al., 2017; Carlini et al., 2022) and reconstruction attacks (e.g. Balle et al., 2022; Nasr et al., 2023). Differential privacy (DP; Dwork et al., 2006b) provides protection against these attacks, but strong formal protection often comes at the cost of significant loss of model utility.
Finding the correct balance between making models resistant to attacks while maintaining a high utility is important for many applications. In health, for example, many European countries and soon also the EU within the European Health Data Space have requirements that models trained on health data that are made publicly available must be anonymous, i.e. they must not contain information that can be linked to an identifiable individual. On the other hand, loss of utility of the model may compromise the health benefits that might be gained from it.
In this paper, our aim is to apply a state-of-the-art MIA systematically to help understand practical privacy risks training deep-learning-based classifiers without DP protections. Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep learning models to MIA. Since the earliest MIA methods, there has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017). Similarly, there is evidence that models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023). Based on extensive experimental data over many data sets with varying sizes, we uncover a power law describing the vulnerability remarkably well.
Previously Tobaben et al. (2023) and Yu et al. (2023) have presented limited studies of how data set properties affect the MIA vulnerability. Tobaben et al. (2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots (i.e. the number of examples available per class). Yu et al. (2023) studied how the MIA vulnerability of a class in an image classifier trained from scratch depends on the average individual privacy parameters of that class. Our work significantly expands on both of these by more systematic study and modelling of data set properties and by providing a more detailed look at individual samples, not just classes.
List of contributions
In this work we perform an extensive empirical study on the MIA vulnerability of deep learning models. We will focus on transfer learning setting on computer vision tasks, where a large pre-trained neural network is fine-tuned on a sensitive data set, and the MIA is performed on the final classifier. We find that the MIA vulnerability has a strong correlation with certain data set properties (Section 5) and make the following contributions:
@@
    Power law: We uncover a power law between the number of examples per class and the vulnerability to MIA (tpr at fixed low fpr) based on extensive experimental data over many data sets with varying sizes (See Figure 1(a)).
@@
    Regression model: We utilize the observations to train a regression model to predict MIA vulnerability (tpr at fixed low fpr) based on examples per class ($S$) and number of classes ($C$) and show both very good fit on the training data as well as good prediction quality on unseen data from a different feature extractor (See Figure 3).
Additionally, in Section 6 we extend the analysis to individual samples and discover a correlation between vulnerability of individual samples and the gradient norms during training. We further study the correlation between sample’s vulnerability and the similarity to the other samples in the feature space.
2 Background
Notation for the properties of the training data set $\mathcal{\mathcal{D}}$:
  • $C$ for the number of classes
  • $S$ for shots (examples per class)
  • $\mathcal{|\mathcal{D}|}$ for training data set size ($\mathcal{|\mathcal{D}|}=CS$)
We denote the number of MIA shadow models with $M$.
@@ 2 Background
LiRA exploits the observation that the value of the loss function used to train a model is often lower for examples that have been part of the training set than for examples that have not been. For a target sample $x$, LiRA trains the shadow models: 1 with $x$ as a part of the training set ($x\in\mathcal{D}$) 2 and without $x$ in the training set ($x\notin\mathcal{D}$) . After training the shadow models, $x$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed, one for the losses of $x\in\mathcal{D}$ shadow models and one for the $x\notin\mathcal{D}$. Finally the attacker computes the loss for the point $x$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $x\in\mathcal{D}$ or $x\notin\mathcal{D}$. Carlini et al. (2022) proposed an optimization for performing LiRA for multiple models and points without training a computationally infeasible amount of shadow models. We utilize this optimization in our experiments.
In Appendix A we provide background on Differential privacy (DP; Dwork et al., 2006b) and DP stochastic gradient descent (DP-SGD) that we use for training DP models.
3 Methods
In this Section, we introduce the preliminaries for evaluating MIA vulnerability.
3.1 Measuring MIA vulnerability
Using the likelihood-ratios from LiRA Carlini et al. (2022) as a score, we can build a binary classifier to predict whether sample belongs to the training data or not. The operational characteristics of such classifier can be used to measure the success of the MIA. More specifically, throughout the rest of the paper, we will use the true positive rate (tpr) at a specific false positive rate (fpr) as a measure for the vulnerability. Identifying even a small number of samples with high confidence is considered harmful (Carlini et al., 2022) and thus we focus on the regions of small fpr.
The tpr and fpr can be also connected to DP. Kairouz et al. (2015) have shown in Equation 1 that any classifier distinguishing the training samples based on the results of a DP algorithm has an upper bound for the tpr that depends on the privacy parameters $(\epsilon,\delta)$. Since MIA is trying to build exactly these types of classifiers, we can use the upper bounds by Kairouz et al. (2015) to validate the privacy claims, and also to better understand the gap between the theoretical privacy guarantees and the realistic attacks.
Theorem 1 (Kairouz et al. (2015)).
A mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-DP if and only if for all adjacent $\mathcal{D}\sim\mathcal{D}^{\prime}$
$\mathrm{TPR}\leq\min\{e^{\epsilon}\mathrm{FPR}+\delta,1-e^{-\epsilon}(1-\delta-\mathrm{FPR})\}\ .$ (1)
3.2 Measuring the uncertainty for TPR
The tpr values from the LiRA based classifier can be seen as maximum likelihood-estimators for the probability of producing true positives among the positive samples. Since we have a finite number of samples for our estimation, it is important to estimate the uncertainty in these estimator. Therefore, when we report the tpr values for single repeat of the learning algorithm, we estimate the stochasticity of the tpr estimate by using Clopper-Pearson intervals Clopper & Pearson (1934). Given tp true positives among p positives, the $1-\alpha$ confidence Clopper-Pearson interval for the tpr is given as
@@ 3.2 Measuring the uncertainty for TPR
4 Experiment overview
Throughout the experiments, we utilize LiRA to understand the vulnerability of the trained models.
We fine-tune pre-trained models on sensitive downstream data sets and assess the vulnerability using LiRA. We use the following setup for all experiments:
  • Pre-trained models. BiT-M-R50x1 (R-50; Kolesnikov et al., 2020) with 23.5M parameters and Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021) with 85.8M parameters, both pre-trained on the ImageNet-21K data set (Russakovsky et al., 2015)
  • Parameterization. Due to computational cost we only consider fine-tuning subsets of the model parameters. We consider the following configurations: 1 Head: training a linear last layer classifier on top of the pre-trained model 2 FiLM: Head and additionally training parameter-efficient FiLM (Perez et al., 2018) adapters scattered throughout the network (see Section B.1)
  • Fine-tuning data sets. We utilize different fine-tuning data sets (see Section B.3 for details) with different $S$, $C$ and $|\mathcal{D}|$ as specified in the experiments.
  • Hyperparameter optimization (HPO). We utilize the library Optuna Akiba et al. (2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations. We optimize the batch size, learning rate and number of epochs. Additionally for DP-SGD we optimize the clipping norm. Details in Section B.2.
  • DP-SGD. We are using Opacus Yousefpour et al. (2021) and compute the noise multiplier based on the target ($\epsilon,\delta$).
  • DP Accounting. We use numerical privacy accounting for accurate evaluation of the cumulative privacy loss of the optimisation (Koskela et al., 2020, 2021) using the PRV accountant (Gopi et al., 2021).
  • LiRA. We base our experiments on the publically available source code of Carlini et al. (2022). We utilize $M=256$ shadow models unless specified otherwise.
We illustrate the effectiveness of LiRA and the effect of different levels of DP on the MIA vulnerability (tpr at $\textsc{fpr}=0.001$) in Figure 1. It shows the MIA vulnerability as a function of privacy budget (additional versions for different fprs in Figure A.1) and two upper bounds on the tpr at fpr emerging from the theoretical guarantees of DP-SGD.
Figure 1: MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of privacy budget ($\epsilon$ at $\delta=10^{-5}$) when attacking different backbones fine-tuned on CIFAR10 and CIFAR100 at $S\in\{25,50\}$. The solid line displays the median and the error bars the minimum of the lower bounds and the maximum of the upper bounds for the Clopper-Pearson CIs over 15 seeds. The upper bound is the theoretical DP bound based on Kairouz et al. (2015). We present the bound computed with $\delta=10^{-5}$ as well as the optimal bound based on multiple $\delta$.
We compute the upper bounds using the method of Kairouz et al. (2015) in Equation 1. Simply using $\delta=10^{-5}$ and the associated $\epsilon$ gives a relatively weak bound, that can be sharpened considerably by finding an optimal $\delta$ and corresponding $\epsilon(\delta)$, which is made possible by the particular mechanism deployed, that gives the tightest possible bound. For the attacked few-shot models the bounds are quite tight for the medium privacy budget of $\epsilon=1$ but very loose at low privacy budgets of $\epsilon=8$ and thus meaningless for low privacy levels.
5 Predicting MIA vulnerability for data sets
@@
In this Section, we investigate how different properties of data sets affect the MIA vulnerability in the non-DP setting. Based on our observations, we propose a method to predict the vulnerability to MIA using these properties of the data set.
We base our experiments on a subset of the few-shot benchmark VTAB  (Zhai et al., 2019) that achieves a classification accuracy $>80\%$ and thus would considered to be used by practitioners. For more detailed descriptions of the data sets, see Table A2. In this Section, we will focus on a fine-tuned last layer classifier (Head) trained on top of a ViT-B, pre-trained on ImageNet-21k Russakovsky et al. (2015). The results for using R-50 as a backbone can be found in Section C.1.
5.1 Correlations with data set properties
@@
Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section 3.1). We make the following observations:
@@
  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1(a). We provide further evidence of this in Figures A.2, A3 and A4 in the Appendix.
@@
(b) as function of $C$ (classes)
Figure 2: MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of dataset properities when attacking a ViT-B Head fine-tuned without DP on different data sets. The solid line displays the median and the error bars the minimum of the lower bounds and maximum of the upper bounds for the Clopper-Pearson CIs over multiple seeds (6 for Figure 1(a) and 12 for Figure 1(b))
@@ 5.2 Model to predict data set vulnerability
Based on the trends seen in Figures 1(b) and 1(a), we fit a linear regression model using the logarithmic transform of $C,S$ to predict the tpr (also logarithmically transformed) with statsmodels Seabold & Perktold (2010). The general form of the model can be found in Equation 3 where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
$\log_{10}(\textsc{tpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (3)
We utilize MIA results of ViT-B (see Table A3) as the training data and use R-50 (see Table A4) as the test data to investigate if the vulnerability prediction model generalizes to a completely different backbone.
@@
The parameters of the prediction model fitted to the training data (note that the parameters are specific to $\textsc{fpr}=0.001$) are $\beta_{S}=-0.492,\beta_{C}=0.357$ and $\beta_{0}=-0.486$. Based on the $R^{2}$ score ($R^{2}=0.917$), our model fits the data extremely well. Furthermore, Figure 3 shows that the model is robust to a change of feature extractor, as it is able to predict the tpr for R-50 despite being trained on ViT-B (test $R^{2}=0.740$). See Section C.1.3 for further supporting results at $\textsc{fpr}\in\{0.1,0.01\}$.
Figure 3: Predicted MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of $S$ (shots) using a model based on Equation 3 fitted Table A3 (ViT-B). The dots show the median tpr for the train set (ViT-B; Table A3) and the test set (R-50; Table A4) over six seeds (datasets: Patch Camelyon, EuroSAT and CIFAR100).

### v2

# On the Impact of Dataset Properties on Membership Privacy of Deep Learning

@@ Abstract
We apply a state-of-the-art membership inference attack (MIA) to systematically test the practical privacy vulnerability of fine-tuning large image classification models. We focus on understanding the properties of data sets and samples that make them vulnerable to membership inference. In terms of data set properties, we find a strong power law dependence between the number of examples per class in the data and the MIA vulnerability, as measured by true positive rate of the attack at a low false positive rate. We train a linear model to predict true positive rate based on data set properties and observe good fit for MIA vulnerability on unseen data. To analyse the phenomenon theoretically, we reproduce the result on a simplified model of membership inference that behaves similarly to our experimental data. We prove that in this model, the logarithm of the difference of true and false positive rates depends linearly on the logarithm of the number of examples per class. For an individual sample, the gradient norm is predictive of its vulnerability.
@@ 1 Introduction
Machine learning models are prone to memorising their training data, which makes them vulnerable to privacy attacks such as membership inference attacks (MIAs; Shokri et al., 2017; Carlini et al., 2022) and reconstruction attacks (e.g. Balle et al., 2022; Nasr et al., 2023). Differential privacy (DP; Dwork et al., 2006) provides protection against these attacks, but strong formal protection often comes at the cost of significant loss of model utility.
Finding the correct balance between making models resistant to attacks while maintaining a high utility is important for many applications. In health, for example, many European countries and soon also the EU within the European Health Data Space have requirements that models trained on health data that are made publicly available must be anonymous, i.e. they must not contain information that can be linked to an identifiable individual. On the other hand, loss of utility of the model due to privacy constraints may compromise the health benefits that might be gained from it.
In this paper, our aim is to apply a state-of-the-art MIA systematically, to help understand practical privacy risks when training deep-learning-based classifiers without DP protections. Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep learning models to MIA. Since the earliest MIA methods, there has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017). Similarly, there is evidence that models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023). Based on extensive experimental data over many data sets with varying sizes, we uncover a power law describing the vulnerability remarkably well.
Previously, Tobaben et al. (2023) and Yu et al. (2023) have presented limited studies of how data set properties affect the MIA vulnerability. Tobaben et al. (2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots (i.e. the number of examples available per class). Yu et al. (2023) studied how the MIA vulnerability of a class in an image classifier trained from scratch depends on the average individual privacy parameters of that class. Our work significantly expands on both of these by more systematic study and modelling of data set properties and by providing a more detailed look at individual samples, not just classes.
List of contributions In this work we perform an extensive empirical study on the MIA vulnerability of deep learning models. We focus on a transfer learning setting for image classification tasks, where a large pre-trained neural network is fine-tuned on a sensitive data set, and the MIA is performed on the final classifier. We find that the MIA vulnerability has a strong correlation with certain data set properties (Section 3). In Section 4 we analyse the reasons for these findings more thoroughly and we extend the analysis to individual samples in Section 5. The primary contributions of this work are:
@@
    Few-shot MIA: We conduct a comprehensive study of MIA vulnerability (tpr at fixed low fpr) of target models trained using many different data sets with varying sizes and uncover a power law between the number of examples per class and the vulnerability to MIA (see Figure 1).
@@
    Regression model: We utilise our empirical observations to train a regression model to predict MIA vulnerability ($\log(\textsc{tpr}-\textsc{fpr})$ at fixed low fpr) based on examples per class ($\log S$) and number of classes ($\log C$). We show both very good fit on the training data as well as good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Figure 3).
  3. 3.
    Theoretical analysis using a simplified model of MIA: To analyse the phenomenon theoretically, we reproduce the result on a simplified model of membership inference that behaves similarly to our experimental data. In this model we can formally prove that $\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log S$ (see Section 4).
  4. 4.
    Individual sample MIA: We extend the study to individual samples and find: 1 The gradient norm is predictive of vulnerability of an individual sample. 2 Gradient norms at different stages of learning have different impact, but how depends on the data set (see Section 5).
2 Background and Methods
Notation for the properties of the training data set $\mathcal{\mathcal{D}}$: 1 $C$for the number of classes 2 $S$for shots (examples per class) 3 $\mathcal{|\mathcal{D}|}$for training data set size ($\mathcal{|\mathcal{D}|}=CS$). We denote the number of MIA shadow models with $M$.
@@ 2 Background and Methods
LiRA exploits the observation that the loss function value used to train a model is often lower for examples that were part of the training set compared to those that were not. For a target sample $x$, LiRA trains the shadow models: 1 with $x$ as a part of the training set ($x\in\mathcal{D}$) 2 and without $x$ in the training set ($x\notin\mathcal{D}$) . After training the shadow models, $x$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed: one for the losses of $x\in\mathcal{D}$ shadow models, and one for the $x\notin\mathcal{D}$. Finally, the attacker computes the loss for the point $x$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $x\in\mathcal{D}$ or $x\notin\mathcal{D}$. Carlini et al. (2022) proposed an optimisation for performing LiRA for multiple models and points without training a computationally infeasible number of shadow models. We utilise this optimisation.
Measuring MIA vulnerability Using the likelihood-ratios from LiRA as a score, we can build a binary classifier to predict whether a sample belongs to the training data or not. The operational characteristics of such classifier can be used to measure the success of the MIA. More specifically, throughout the rest of the paper, we will use the true positive rate (tpr) at a specific false positive rate (fpr) as a measure for the vulnerability. Identifying even a small number of samples with high confidence is considered harmful (Carlini et al., 2022) and thus we focus on the regions of small fpr.
Measuring the uncertainty for tpr The tpr values from the LiRA-based classifier can be seen as maximum likelihood-estimators for the probability of producing true positives among the positive samples. Since we have a finite number of samples for our estimation, it is important to estimate the uncertainty in these estimators. Therefore, when we report the tpr values for a single repeat of the learning algorithm, we estimate the stochasticity of the tpr estimate by using Clopper-Pearson intervals Clopper and Pearson (1934). Given tp true positives among p positives, the $1-\alpha$ confidence Clopper-Pearson interval for the tpr is given as
@@
3 MIA Vulnerability and Data Set Properties
@@
In this section, we investigate how different properties of data sets affect the MIA vulnerability. Based on our observations, we propose a method to predict the vulnerability to MIA using these properties.
3.1 Experimental Setup
We focus on a image classification setting where we fine-tune pre-trained models on sensitive downstream data sets and assess the MIA vulnerability using LiRA with $M=256$ shadow models. We base our experiments on a subset of the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a test classification accuracy $>80\%$ (see Table A2).
We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k Russakovsky et al. (2015). The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section B.1. We optimise the hyperparameters (batch size, learning rate and number of epochs) using the library Optuna Akiba et al. (2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section A.2). We provide the the code for reproducing the experiments in the supplementary material.
3.2 Experimental Results
@@
Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section 2). We make the following observations:
@@
  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1(a). We provide further evidence of this in the Appendix (Figures A.1, A3 and A4).
@@
(b) as function of $C$ (classes) at $S=32$
Figure 1: MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of data set properities when attacking a ViT-B Head fine-tuned without DP on different data sets. The solid line displays the median and the error bars the minimum of the lower bounds and maximum of the upper bounds for the Clopper-Pearson CIs over multiple seeds (six for Figure 1(a) and 12 for Figure 1(b))
@@ 3.3 Model to Predict Data Set Vulnerability
[…]
$\log_{10}(\textsc{tpr}-\textsc{fpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (2)
[…]
@@
We utisee MIA results of ViT-B (see Table A3) as the training data. Based on the $R^{2}$ (coefficient of determination) score ($R^{2}=0.930$ for the model trained on $\textsc{fpr}=0.001$ data), our model fits the data extremely well. We provide further evidence for other fpr in Figure A.3 and Table A8 in the Appendix. Figure 2 shows the parameters of the prediction model fitted to the training data.
Figure 2: Coefficient values for different fpr when fitting a regression model based on Equation 2 fitted on data from ViT-B (Head) (Table A3). The error bars display the 95% confidence intervals based on Student’s t-distribution.
Prediction quality on other MIA target models
We analyse how the regression model trained on the ViT-B (Head) data generalizes to other target models. The main points are:
  • R-50 (Head): Figure 3(a) shows that the regression model is robust to a change of the feature extractor, as it is able to predict the tpr for R-50 (Head) (test $R^{2}=0.790$).
  • R-50 (FiLM): Figure 3(b) shows that the prediction quality is good for R-50 (FiLM) models. These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section A.1). Tobaben et al. (2023) demonstrated that FiLM layers are a competitive alternative to training all parameters. We supplement the MIA results of Tobaben et al. (2023) with own FiLM training runs. Refer to Table A7 in the Appendix.
  • From-Scratch-Training: Carlini et al. (2022) provide limited results on from-scratch-training. To the best of our knowledge these are the only published LiRA results on image classification models. Figure 3(b) displays that our prediction model underestimates the vulnerability of the from-scratch trained target models. We have identified two potential explanations for this 1 In from-scratch-training all weights of the model need to be trained from the sensitive data and thus potentially from-scratch-training could be more vulnerable than fine-tuning. 2 The strongest attack in Carlini et al. (2022) uses data augmentations to improve the performance. We are not using this optimization.
[…]
(b) Evaluating prediction performance for target models 1 left: fine-tuned with FiLM (data partially from Tobaben et al. (2023), see Table A7) 2 right: trained from scratch (data from Carlini et al. (2022))
Figure 3: Performance of the regression model based on Equation 2 fitted on data from Table A3.
4 Understanding the MIA Prediction Model
In this section, we focus on understanding why the previously observed trends exist. Our goal is to formally prove a similar relationship as in the predictive model of Section 3.3.
4.1 A Simplified Model of Membership Inference
In this section, we construct a simplified model of membership inference that streamlines the data generation and shadow model training. We observe that the trends of the vulnerability of the simplified model is a good approximation of the trends previously observed with the image classification example.
Instead of training linear classifiers to classify images, we sample vectors on a high dimensional unit sphere and classify based on inner product with estimated class mean. This model is easier to analyse theoretically, and allows us to simulate arbitrary combinations of $S$ and $C$ without relying on the existence of a specific image classification downstream data set as especially high $C$ image classification downstream data sets that are computationally feasible to fine-tune on have small $S$. We generate the data and form the classifiers (which are our target models) as follows:
  1. 1.
    For each class, we first sample a true class mean $m_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in C:m_{i}\perp m_{j}\lor i=j$).
  2. 2.
    We sample $2S$ vectors $x_{c}$ for each class. We assume that they are Gaussian distributed around the the true class mean $x_{c}\sim\mathcal{N}(m_{c},s^{2}I)$ where the $s^{2}$ is the in-class variance.
  3. 3.
    For each "target model" we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $r_{c}$.
  4. 4.
    The computed mean is used to classify sample $(x,y)$ by computing the inner product $\langle x,r_{c}\rangle$ as a metric of similarity.
The attacker has to infer which vectors have been used for training the classifier. Instead of utilising the logits (like in the image classification example), the attacker can use the inner products of a point with the cluster means. In Figure 4 we compare the attacker’s performance as a function of $C$ and $S$ to the the previous image classification example.
This simplified model resembles the linear (Head) classifier. In the linear classifier, we find a matrix $A$ and biases $b$, to optimize the cross-entropy between the labels and logits $Az+b$, where $z$ denotes the feature space representation of the data. In the simplified model, the rows of $A$ are replaced by the cluster means and we do not include the bias term in the classification.
(a) as a function of $S$ (shots)
(b) as a function of $C$ (classes)
[…]
4.2 Theoretical Analysis
Similarly to the setting in Section 4.1, we focus on a classification model that aims to classify the sample $(x,y)$ based on the inner product $\langle x,r_{c}\rangle$, where $r_{c}$ is the mean of the training data points $x_{m}$ with $y_{m}=c$. However, we only focus on the scores of the target class, i.e. $c=y$, and therefore the score used of LiRA simplifies to $\langle x,r_{y}\rangle$.
The fpr for LiRA is given as
$\textsc{fpr}=\Pr\left(\frac{\mathcal{N}(t_{x};\mu_{\textrm{in}},\sigma_{\textrm{in}}^{2})}{\mathcal{N}(t_{x};\mu_{\textrm{out}},\sigma_{\textrm{out}}^{2})}>\exp(\tau)\mid x\notin\mathcal{D}_{\textrm{target}}\right),$ (3)
where $t_{x}=\langle x,r_{y}^{\textrm{(target)}}\rangle$ is the score for the target model, $\tau$ is the decision threshold for the classifier, and $\mu_{\textrm{in}},\mu_{\textrm{out}},\sigma_{\textrm{in}}$ and $\sigma_{\textrm{out}}$ are the LiRA statistics computed from the shadow models. We assume that $\sigma_{\textrm{in}}=\sigma_{\textrm{out}}=\sigma$, which allows to write Equation 3 as
fpr $=\Pr\left(\log\frac{\mathcal{N}(t_{x};\mu_{\textrm{in}},\sigma^{2})}{\mathcal{N}(t_{x};\mu_{\textrm{out}},\sigma^{2})}>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right)$ (4)
$=\Pr\left(-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right).$ (5)
From now on, we will denote the $-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]$ with $\lambda_{\textsc{fpr}}$ when $x\notin\mathcal{D}_{\textrm{target}}$ and with $\lambda_{\textsc{tpr}}$ otherwise. Furthermore we will drop the condition $x\in\mathcal{D}_{\textrm{target}}$ or $x\notin\mathcal{D}_{\textrm{target}}$ from the fpr and tpr expressions
fpr $:=\Pr(\lambda_{\textsc{fpr}}>\tau)$ (6)
tpr $:=\Pr(\lambda_{\textsc{tpr}}>\tau).$ (7)
Lemma 1.
In the limit of infinite shadow models and with large enough $S$, the LiRA test statistics $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ obtain the following forms
$\lambda_{\textsc{fpr}}$ $=-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle+\frac{1}{S}\langle x,x-m_{y}\rangle\right)$ (8)
$\lambda_{\textsc{tpr}}$ $\approx-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle-\frac{2}{S}||x||^{2}+\frac{1}{S}\langle x,x-m_{y}\rangle\right),$ (9)
where $z\sim\mathcal{N}(0,I_{d})$.
Proof.
We defer the proof to Appendix C. ∎
The factors independent of $S$ in for both $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ correspond to the $-(\mu_{\textrm{in}}-\mu_{\textrm{out}})/(2\sigma^{2})$. Figure A.5 illustrates how this factor behaves as a function of $S$ in our experiments with the simple model. As predicted by the theory, the factor (i.e. the difference in means divided by the variance) remains fairly constant across different values of $S$.
Using Lemma 1 we can derive a rate for the decision threshold $\tau$:
Corollary 2.
In order to maintain constant fpr, we need to adjust $\tau$ as $O(1/\sqrt{S})$.
According to Corollary 2, the $\tau:=\tau(S)$ evolves as $O(1/\sqrt{S})$, and we will approximate it as $\tau(S)=\gamma_{1}\frac{1}{\sqrt{S}}$, with constant $\gamma_{1}$ depending on the target fpr.
Theorem 3.
$\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log(S)$.
Proof.
From Lemma 1, we see that $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ differ only in a single term which evolves as $O(1/S)$. Thus for sufficiently large $S$ we can consider this additional r.v. as a small constant shift to the distribution of $\lambda_{\textsc{fpr}}$. Let us denote this variable now with $q$. When $S$ grows, $q$ becomes small, and we have
[…]
$\approx p_{\lambda_{\textsc{fpr}}}(\tau)q.$ (11)
Taking the log yields
$\log(\textsc{tpr}-\textsc{fpr})$ $\approx\log\left(p_{\lambda_{\textsc{fpr}}}\left(\tau\right)\right)+\log q$ (12)
$=\log\left(p_{\lambda_{\textsc{fpr}}}\left(\tau\right)\right)-\log S+\gamma_{2}.$ (13)
From Lemma 1, we can see that a scaled r.v. $\hat{\lambda}_{\textsc{fpr}}:=\sqrt{S}\lambda_{\textsc{fpr}}$ becomes almost independent of $S$ for large enough $S$. Hence using the change of variables formula, we have that $p_{\lambda_{\textsc{fpr}}}(t)=p_{\hat{\lambda}_{\textsc{fpr}}}(\sqrt{S}t)\sqrt{S}$. Now, evaluating $p_{\lambda_{\textsc{fpr}}}(t)$ at $t=\tau(S)$, we get $p_{\lambda_{\textsc{fpr}}}(t)=p_{\hat{\lambda}_{\textsc{fpr}}}(\gamma_{1})\sqrt{S}$, where the $p_{\hat{\lambda}_{\textsc{fpr}}}(\gamma_{1})$ is independent of $S$. Therefore, ignoring additive constants,
$\log(\textsc{tpr}-\textsc{fpr})$ $\approx\log\sqrt{S}-\log S=\frac{1}{2}\log S-\log S=-\frac{1}{2}\log S.$ (14)
This shows that $\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log S$. ∎
Compared to the model in Section 3.3, our theoretical model predicts $\beta_{S}=-1/2$. Our empirical estimates match this for high fpr. For lower FPR, $\beta_{S}$ decreases, which indicates there may be finer details of $p_{\lambda_{\textsc{fpr}}}$ we are missing in this analysis.